A thin uniform circular strip, flexural rigidity EI, fixed at C and is free to slide without friction at A. The strip carries a vertical force F at B as shown. Determine the horizontal reaction at A using Castigliano’s theorem. Answer.411F

B TTT ाIो \45 F
Answer #1


Let Horisontal reed at A2 HA Bl mus Bi AB part 1. PR-rcose) - HA Ma = HA *(R-R cose) Streen energy in AB U=(M² R de ZEI 23aly de + S [ Halltsno) – Fli-caso) ] * Clesino) do 12 7.0 +sir 20 T x 2) 0 = Ha Te - 2 bio 2. Jo [ac-po-200 (9 ). to co - 21. Oz Hay ( 311 - 2) + HAX (7-198)-fi (3-1062) 2) Haz 3-1062f 7.1qg to. 3562 Hac 0.411f